OpenStudy (anonymous):

What is the probability that out of a group of 2 people, both will have same birthday. Assuming that there are 365 days and none have their birthday on 29th February.

OpenStudy (anonymous):

@Astrophysics could you please help?

OpenStudy (anonymous):

You can use binomial expansion. n=365 p=1/365 (probability of success of a trial) q=1-p=1-1/365=364/365 (probability of failure of a trial) The original formula is given by\[\left( nCx \right) *p ^{x}*q ^{n-x}\] Where x is the number of desired successful trials. Probability will than be P(X=2)\[(365C2)*(\frac{ 1 }{ 365 })^{2}*(\frac{ 364 }{ 365 })^{365-2}=0.184\]

OpenStudy (anonymous):

@Astrophysics